# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ V Conclusions
where $\sigma_{k-1}=\delta_{k-1}\gamma_{k-1}$, and we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\gamma_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$. However, when no feedback information is delivered in the backward channel at time $k$, we have $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$. Taking the conditional expectation and the conditional covariance of (29) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$, i.e., when $\delta_{k-1}=0$ or $\delta_{k-1}=1\wedge\theta_{k-1}=0$, respectively, we get
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}+K_{k}\nu_{k}$ (34)
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\lambda^{c}\delta_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}A_{k-1}^{T}$
$\quad+(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})A_{k-1}$
×$\quad\times\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\Big{|}\mathcal{I}^{e}_{k-1}\Big{]}^{T}A_{k-1}^{T}$ (35)
@@
where we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{G}\}$ or $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{H}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\gamma_{k-1}$ is independent of $\tilde{e}_{k-1}$ and $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$, and the properties of covariances. Now, if $\mu_{k-1}\neq 0$, it is easy to see that the contributions of $\mu_{k-1}$ in $\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ are captured by $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$ and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]+2\operatorname{\mathsf{cov}}[\mu_{k-1},(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$, respectively. As a result, following the definitions of $\breve{e}_{k}$ and $R_{k}$, we can obtain (30) by combining $\varsigma_{k-1}$ with (32) and (34), and can obtain (31) by combining $\Xi_{k-1}$ with (33) and (35). ∎
@@
Let $\operatorname{\mathsf{p}}(\delta_{k}=0|\boldsymbol{\nu}_{0:k},\boldsymbol{\delta}_{0:k-1})$ be a symmetric function of $\boldsymbol{\nu}_{0:k}$ for $k\in\mathbb{N}_{[0,N]}$. Then, $\imath_{k}$, $\jmath_{k}$, $\mu_{k}$, $\varsigma_{k}$, and $\Xi_{k}$ in Lemmas 3 and 4, are equal to zero for $k\in\mathbb{N}_{[0,N-1]}$.
@@
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\delta}_{0:t}\big{|}\boldsymbol{\nu}_{0:t}\Big{)}=\prod_{t^{\prime}=0}^{t}\operatorname{\mathsf{p}}\Big{(}\delta_{t^{\prime}}\big{|}\boldsymbol{\nu}_{0:t^{\prime}},\boldsymbol{\delta}_{0:t^{\prime}-1}\Big{)}$
where we used the fact that $\delta_{t^{\prime}}$ is independent of $\boldsymbol{\nu}_{t^{\prime}+1:t}$ given $\boldsymbol{\nu}_{0:t^{\prime}}$ and $\boldsymbol{\delta}_{0:t^{\prime}-1}$. Therefore, by the hypothesis, we can deduce that $\operatorname{\mathsf{p}}(\boldsymbol{\delta}_{0:t}|\boldsymbol{\nu}_{0:t})$ is symmetric with respect to $\boldsymbol{\nu}_{0:t}$. It follows that
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{0:t}\big{|}\boldsymbol{\delta}_{0:t}\Big{)}\propto\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\delta}_{0:t}\big{|}\boldsymbol{\nu}_{0:t}\Big{)}\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{0:t}\Big{)}.$
Since $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t})$ is a symmetric distribution, we deduce that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t}|\boldsymbol{\delta}_{0:t})$ is also symmetric with respect to $\boldsymbol{\nu}_{0:t}$. For any $s\in\mathbb{N}_{[0,t]}$, by marginalization and the fact that $\boldsymbol{\nu}_{s:t}$ is independent of $\boldsymbol{\delta}_{0:s-1}$, we can find that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{s:t}|\boldsymbol{\delta}_{s:t})$ is symmetric with respect to $\boldsymbol{\nu}_{s:t}$. Furthermore, for any $k\in\mathbb{N}_{[t+1,N]}$, we have
$\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{s:k}\big{|}\boldsymbol{\delta}_{s:t}\Big{)}=\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{s:t}\big{|}\boldsymbol{\delta}_{s:t}\Big{)}\operatorname{\mathsf{p}}\Big{(}\boldsymbol{\nu}_{t+1:k}\Big{)}.$
Therefore, we can see that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{s:k}|\boldsymbol{\delta}_{s:t})$ is also symmetric with respect to $\boldsymbol{\nu}_{s:k}$.
Note that $z^{f}_{k+1}=\mathfrak{E}$ and $z^{f}_{k+1}=\mathfrak{F}$ are equivalent to $\delta_{k}=0$ and $\{\delta_{k}=1,\gamma_{k}=0\}$, respectively. Accordingly, when $z^{f}_{k+1}=\mathfrak{E}$ or $z^{f}_{k+1}=\mathfrak{F}$, we can write
@@
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k},z^{f}_{k+1}\Big{]}\Big{|}\mathcal{I}^{d}_{k},z_{k+1}\bigg{]}$
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}\Big{|}\mathcal{I}^{d}_{k},z_{k+1}\bigg{]}$
$\qquad=\operatorname{\mathsf{E}}\bigg{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg{]}$
where in the first equality we used the tower property of conditional expectations, and in the second equality the facts that $\hat{e}_{k}$ is independent of $\gamma_{k}$, and that $\delta_{k}\in\mathcal{I}^{e}_{k}$. In addition, from (29), when $\imath_{t}=\jmath_{t}=0$ for all $t\in\mathbb{N}_{[d-1,k-1]}$, we can write
@@
Let us define $a_{t},b_{t}\in\{0,1\}$ such that $a_{t}b_{t}=0$ for $t\in\mathbb{N}_{[0,N]}$. In the case of the channel with packet-loss detection, by (Proof:) and the definitions of $\imath_{k}$ and $\jmath_{k}$, we obtain
@@
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{1}_{k+1}\bigg{]}$
@@
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\mathcal{I}^{2}_{k+1}\bigg{]}$
where $\mathcal{I}^{1}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=0,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$ and $\mathcal{I}^{2}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=1,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\gamma_{k}$. Equivalently, we can write
$\imath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\bar{\mathcal{I}}^{1}_{k+1}\bigg{]}$
$\jmath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg{[}\boldsymbol{\nu}_{k-\eta_{k}:k}\Big{|}\bar{\mathcal{I}}^{2}_{k+1}\bigg{]}$
where ¯$\bar{\mathcal{I}}^{1}_{k+1}=\{\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=0\}$ and ¯$\bar{\mathcal{I}}^{2}_{k+1}=\{\boldsymbol{\delta}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},\delta_{k}=1\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\mathcal{I}^{d}_{k-\eta_{k}}$ and $\boldsymbol{\gamma}_{k-\eta_{k}:k-1}$. We have already shown that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{k-\eta_{k}:k}|\boldsymbol{\delta}_{k-\eta_{k}:k})$ is symmetric with respect to $\boldsymbol{\nu}_{k-\eta_{k}-d+1:k}$. This implies that $\imath_{k}=0$ and $\jmath_{k}=0$. ∎
The next two technical lemmas are related to symmetric decreasing rearrangements of non-negative functions. For the proofs of these lemmas, see, e.g., [58] and [59].
@@
[…]
@@
Let $\mathcal{B}(r)\subseteq\mathbb{R}^{n}$ be a ball of radius $r$ centered at the origin, and $f$ and $g$ be non-negative functions defined on $\mathbb{R}^{n}$ that vanish at infinity and satisfy
[…]
@@
The next lemma provides an equivalent loss function.
@@
Let $\tilde{e}_{0}$ satisfies $K_{0}\nu_{0}$ and $\delta$ be constructed based on $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. Optimizing the loss function $\Phi(\epsilon,\delta)$ is equivalent to optimizing the following loss function $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ over $\epsilon\in\mathcal{E}$:
@@
×$\times\operatorname{\mathsf{E}}_{\epsilon}\Big{[}\delta_{k}\big{|}\boldsymbol{\sigma}_{0:k-1}=0\Big{]}$
@@
×$\times\operatorname{\mathsf{E}}_{\epsilon}\Big{[}\Psi^{k+1,N}_{\epsilon}(\tilde{e}_{k+1})\big{|}\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big{]}\Big{\}}$ (40)
where $\sigma_{k}=\delta_{k}\gamma_{k}$ and
$\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})=\sum_{t=k}^{N}\operatorname{\mathsf{E}}\Big{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{]}$
for $k\in\mathbb{N}_{[1,N]}$ when $\tilde{e}_{k}$ satisfies $K_{k}\nu_{k}$.
@@
$\Phi(\epsilon,\delta)$ $=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\bigg{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\operatorname{\mathsf{E}}\Big{[}\alpha_{k}\delta_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\big{|}\mathcal{I}^{e}_{k}\Big{]}\bigg{]}$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}+\operatorname{tr}Q_{k}\bigg{]}$ (41)
@@
$\operatorname{\mathsf{E}}\bigg{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{|}\sigma_{k}=1\bigg{]}$
$\quad=\operatorname{\mathsf{E}}\bigg{[}\alpha_{t}\delta_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big{|}\boldsymbol{\sigma}_{0:k-d},\sigma_{k}=1\bigg{]}$ (42)
for any $t\in\mathbb{N}_{[k+1,N]}$, where we used the fact that, without loss of optimality, $\boldsymbol{\delta}_{k+1:N}$ and $\tilde{\boldsymbol{e}}_{k+1:N}$ are independent of $\boldsymbol{\sigma}_{0:k-1}$ given $\sigma_{k}=1$, i.e., when $\tilde{e}_{k+1}=K_{k+1}\nu_{k+1}$.
@@
$\Psi^{N}_{\epsilon}(\tilde{e}_{0}):=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg{[}\alpha_{k}\delta_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}\bigg{]}.$ (43)
Following the facts that $\tilde{e}_{0}$ and $\operatorname{tr}Q_{t}$ for $t\in\mathbb{N}_{[0,N]}$ are independent of the scheduling policy, to optimize $\Phi(\epsilon,\delta)$ in (Proof:) over $\epsilon\in\mathcal{E}$, it suffices to optimize $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ in (43) over $\epsilon\in\mathcal{E}$. In addition, observe that from the law of total probability, we have

### v2

@@ VIII Conclusions
where $\sigma_{k-1}=u_{k-1}\gamma_{k-1}$, and we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\gamma_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$. However, when no feedback information is delivered in the backward channel at time $k$, we have $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$. Taking the conditional expectation and the conditional covariance of (67) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\mathfrak{G}$ or $z^{b}_{k}=\mathfrak{H}$, i.e., $u_{k-1}=0$ or $u_{k-1}=1\wedge\theta_{k-1}=0$, respectively, we get
$\operatorname{\mathsf{E}}\Big[\tilde{e}_{k}\big|\mathcal{I}^{e}_{k}\Big]$ $=(1\!-\!\lambda^{c}u_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big[\tilde{e}_{k-1}\big|\mathcal{I}^{e}_{k-1}\Big]$
$\ \ +K_{k}\nu_{k},$ (72)
$\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k}\big|\mathcal{I}^{e}_{k}\Big]$ $=(1-\lambda^{c}u_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k-1}\big|\mathcal{I}^{e}_{k-1}\Big]A_{k-1}^{T}$
$\ \ +(\lambda^{c}u_{k-1}-{\lambda^{c}}^{2}u_{k-1})A_{k-1}$
$\ \ \times\operatorname{\mathsf{E}}\Big[\tilde{e}_{k-1}\big|\mathcal{I}^{e}_{k-1}\Big]\operatorname{\mathsf{E}}\Big[\tilde{e}_{k-1}\Big|\mathcal{I}^{e}_{k-1}\Big]^{T}A_{k-1}^{T},$ (73)
@@
where we used the facts that $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{G}\}$ or $\{\mathcal{I}^{e}_{k-1},y_{k},z^{b}_{k}=\mathfrak{H}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$, that $\gamma_{k-1}$ is independent of $\tilde{e}_{k-1}$ and $\mathcal{I}^{e}_{k}$, that $\tilde{e}_{k-1}$ is independent of $y_{k}$ and $z^{b}_{k}$ given $\mathcal{I}^{e}_{k-1}$, and that $\nu_{k}$ is known given $\mathcal{I}^{e}_{k}$, and the properties of covariances. Now, if $\mu_{k-1}\neq 0$, which happens only if $u_{k-1}\gamma_{k-1}=0$, it is easy to see that the contributions of $\mu_{k-1}$ in $\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$ are captured by $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]$ and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]+2\operatorname{\mathsf{cov}}[\mu_{k-1},A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]$, respectively. As a result, following the definitions of $\breve{e}_{k}$ and $R_{k}$, we can obtain (68) by combining (70) and (72) and incorporating $\varsigma_{k-1}$, and can obtain (69) by combining (71) and (73) and incorporating $\Xi_{k-1}$. ∎
@@
Let $\operatorname{\mathsf{p}}(u_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{u}_{0:k-1})$ be a symmetric function of $\boldsymbol{\nu}_{0:k}$ for $k\in\mathbb{N}_{[0,N]}$. Then, $\imath_{k}$, $\jmath_{k}$, $\mu_{k}$, $\varsigma_{k}$, and $\Xi_{k}$ in Lemmas 3 and 4, are equal to zero for $k\in\mathbb{N}_{[0,N-1]}$.
@@
$\operatorname{\mathsf{p}}\Big(\boldsymbol{u}_{0:t}\big|\boldsymbol{\nu}_{0:t}\Big)=\prod_{t^{\prime}=0}^{t}\operatorname{\mathsf{p}}\Big(u_{t^{\prime}}\big|\boldsymbol{\nu}_{0:t^{\prime}},\boldsymbol{u}_{0:t^{\prime}-1}\Big),$
where we used the fact that $u_{t^{\prime}}$ is independent of $\boldsymbol{\nu}_{t^{\prime}+1:t}$ given $\boldsymbol{\nu}_{0:t^{\prime}}$ and $\boldsymbol{u}_{0:t^{\prime}-1}$. Therefore, by the hypothesis, we can deduce that $\operatorname{\mathsf{p}}(\boldsymbol{u}_{0:t}|\boldsymbol{\nu}_{0:t})$ is symmetric with respect to $\boldsymbol{\nu}_{0:t}$. It follows that
$\operatorname{\mathsf{p}}\Big(\boldsymbol{\nu}_{0:t}\big|\boldsymbol{u}_{0:t}\Big)\propto\operatorname{\mathsf{p}}\Big(\boldsymbol{u}_{0:t}\big|\boldsymbol{\nu}_{0:t}\Big)\operatorname{\mathsf{p}}\Big(\boldsymbol{\nu}_{0:t}\Big).$
Since $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t})$ is a symmetric distribution, we deduce that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{0:t}|\boldsymbol{u}_{0:t})$ is also symmetric with respect to $\boldsymbol{\nu}_{0:t}$. For any $s\in\mathbb{N}_{[0,t]}$, by marginalisation and the fact that $\boldsymbol{\nu}_{s:t}$ is independent of $\boldsymbol{u}_{0:s-1}$, we can find that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{s:t}|\boldsymbol{u}_{s:t})$ is symmetric with respect to $\boldsymbol{\nu}_{s:t}$.
Note that $z^{f}_{k+1}=\mathfrak{E}$ and $z^{f}_{k+1}=\mathfrak{F}$ are equivalent to $u_{k}=0$ and $\{u_{k}=1,\gamma_{k}=0\}$, respectively. Accordingly, when $z^{f}_{k+1}=\mathfrak{E}$ or $z^{f}_{k+1}=\mathfrak{F}$, we can write
@@
$\qquad=\operatorname{\mathsf{E}}\bigg[\operatorname{\mathsf{E}}\Big[\hat{e}_{k}\big|\mathcal{I}^{e}_{k},z^{f}_{k+1},\boldsymbol{\gamma}_{0:k-1}\Big]\Big|\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg]$
$\qquad=\operatorname{\mathsf{E}}\bigg[\operatorname{\mathsf{E}}\Big[\hat{e}_{k}\big|\mathcal{I}^{e}_{k},\boldsymbol{\gamma}_{0:k-1}\Big]\Big|\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg]$
$\qquad=\operatorname{\mathsf{E}}\bigg[\tilde{e}_{k}\Big|\mathcal{I}^{d}_{k},z^{f}_{k+1}\bigg],$
where in the first equality we used the tower property of conditional expectations, and in the second equality the facts that $\hat{e}_{k}$ is independent of $\gamma_{k}$, and that $u_{k}\in\mathcal{I}^{e}_{k}$. In addition, from (67), when $\imath_{t}=\jmath_{t}=0$ for all $t\in\mathbb{N}_{[d-1,k-1]}$, we can write
@@
Let us define $a_{t},b_{t}\in\{0,1\}$ such that $a_{t}b_{t}=0$ for $t\in\mathbb{N}_{[0,N]}$. By (Proof:) and the definitions of $\imath_{k}$ and $\jmath_{k}$, we obtain
@@
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg[\boldsymbol{\nu}_{k-\eta_{k}:k}\Big|\mathcal{I}^{1}_{k+1}\bigg],$
@@
$=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg[\boldsymbol{\nu}_{k-\eta_{k}:k}\Big|\mathcal{I}^{2}_{k+1}\bigg],$
where $\mathcal{I}^{1}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{u}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},u_{k}=0,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$ and $\mathcal{I}^{2}_{k+1}=\{\mathcal{I}^{d}_{k-\eta_{k}},\boldsymbol{u}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},u_{k}=1,\boldsymbol{\gamma}_{k-\eta_{k}:k-1}=\boldsymbol{b}_{k-\eta_{k}:k-1}\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\gamma_{k}$. Equivalently, we can write
$\imath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg[\boldsymbol{\nu}_{k-\eta_{k}:k}\Big|\bar{\mathcal{I}}^{1}_{k+1}\bigg],$
$\jmath_{k}$ $=A_{k}D_{k}\operatorname{\mathsf{E}}\bigg[\boldsymbol{\nu}_{k-\eta_{k}:k}\Big|\bar{\mathcal{I}}^{2}_{k+1}\bigg],$
where $\bar{\mathcal{I}}^{1}_{k+1}=\{\boldsymbol{u}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},u_{k}=0\}$ and $\bar{\mathcal{I}}^{2}_{k+1}=\{\boldsymbol{u}_{k-\eta_{k}:k-1}=\boldsymbol{a}_{k-\eta_{k}:k-1},u_{k}=1\}$, and we used the fact that $\boldsymbol{\nu}_{k-\eta_{k}:k}$ is independent of $\mathcal{I}^{d}_{k-\eta_{k}}$ and $\boldsymbol{\gamma}_{k-\eta_{k}:k-1}$. We have already shown that $\operatorname{\mathsf{p}}(\boldsymbol{\nu}_{k-\eta_{k}:k}|\boldsymbol{u}_{k-\eta_{k}:k})$ is symmetric with respect to $\boldsymbol{\nu}_{k-\eta_{k}-d+1:k}$. This implies that $\imath_{k}=0$ and $\jmath_{k}=0$, and consequently completes the proof. ∎
The next two technical lemmas are related to the symmetric decreasing rearrangements of non-negative functions. For the proofs of these lemmas, see, e.g., [76] and [77].
@@
$\int_{\mathbb{R}^{n}}f(x)g(x)dx\leq\int_{\mathbb{R}^{n}}f^{*}(x)g^{*}(x)dx,$ (75)
where $f^{*}(x)$ and $g^{*}(x)$ are the symmetric decreasing rearrangements of $f(x)$ and $g(x)$.
@@
Let $\mathcal{B}(r)\subseteq\mathbb{R}^{n}$ be a ball of radius $r$ centred at the origin, and $f$ and $g$ be non-negative functions defined on $\mathbb{R}^{n}$ that vanish at infinity and satisfy
$\int_{\mathcal{B}(r)}f^{*}(x)dx\leq\int_{\mathcal{B}(r)}g^{*}(x)dx,$ (76)
@@
The next lemma provides a loss function that is equivalent to the one under study.
@@
Let $\tilde{e}_{0}$ be equal to $K_{0}\nu_{0}$ and $\delta$ be constructed based on $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. Then, optimising the loss function $\Phi(\epsilon,\delta)$ over $\epsilon\in\mathcal{E}$ is equivalent to optimising the loss function
@@
$\times\operatorname{\mathsf{E}}_{\epsilon}\Big[u_{k}\big|\boldsymbol{\sigma}_{0:k-1}=0\Big]$
@@
$\times\operatorname{\mathsf{E}}_{\epsilon}\Big[\Psi^{k+1,N}_{\epsilon}(\tilde{e}_{k+1})\big|\boldsymbol{\sigma}_{0:k-1}=0,\sigma_{k}=1\Big]\Big\},$ (78)
over $\epsilon\in\mathcal{E}$, where $\sigma_{k}=u_{k}\gamma_{k}$ and
$\Psi^{k,N}_{\epsilon}(\tilde{e}_{k})=\sum_{t=k}^{N}\operatorname{\mathsf{E}}\Big[\alpha_{t}u_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big],$
for $k\in\mathbb{N}_{[1,N]}$ given $\tilde{e}_{k}=K_{k}\nu_{k}$.
@@
$\Phi(\epsilon,\delta)$ $=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg[\alpha_{k}u_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\bigg]$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg[\operatorname{\mathsf{E}}\Big[\alpha_{k}u_{k}+\hat{e}_{k}^{T}\hat{e}_{k}\big|\mathcal{I}^{e}_{k},\boldsymbol{\gamma}_{0:k-1}\Big]\bigg]$
$=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg[\alpha_{k}u_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}+\operatorname{tr}Q_{k}\bigg],$ (79)
@@
$\operatorname{\mathsf{E}}\bigg[\alpha_{t}u_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big|\sigma_{k}=1\bigg]$
$\quad=\operatorname{\mathsf{E}}\bigg[\alpha_{t}u_{t}+\tilde{e}_{t}^{T}\tilde{e}_{t}\Big|\boldsymbol{\sigma}_{0:k-1},\sigma_{k}=1\bigg],$ (80)
for any $t\in\mathbb{N}_{[k+1,N]}$, where we used the fact that, without loss of optimality, $\boldsymbol{u}_{k+1:N}$ and $\tilde{\boldsymbol{e}}_{k+1:N}$ are independent of $\boldsymbol{\sigma}_{0:k-1}$ given $\tilde{e}_{k+1}=K_{k+1}\nu_{k+1}$.
@@
$\Psi^{N}_{\epsilon}(\tilde{e}_{0}):=\sum_{k=0}^{N}\operatorname{\mathsf{E}}\bigg[\alpha_{k}u_{k}+\tilde{e}_{k}^{T}\tilde{e}_{k}\bigg].$ (81)
Following the facts that $\tilde{e}_{0}$ and $\operatorname{tr}Q_{t}$ for $t\in\mathbb{N}_{[0,N]}$ are independent of the scheduling policy, to optimise $\Phi(\epsilon,\delta)$ in (Proof:) over $\epsilon\in\mathcal{E}$, it suffices to optimise $\Psi^{N}_{\epsilon}(\tilde{e}_{0})$ in (81) over $\epsilon\in\mathcal{E}$. In addition, observe that from the law of total probability, we have